OpenStudy (anonymous):

If A= log(base2)2 and B= log(base2)3 write the following in terms of A and B. a) log(base2) (20.25)

OpenStudy (anonymous):

log20.25=log4.5.5.5 log2^2+log5^3 2log2+3log5 2log2=2A 3log(5+1-1)=3(log6-log1) adding 1 and subtracting 1 wont change the solution. 3(log2.3-log1) 3(log2+log3-log1) log1=0 3(log2+log3)=3A+3B 2A+3A+3B=5A+3B ididnt type the bases all bases are 2 remember the log rules loga.b=loga+logb log(a/b)=loga-logb loga^n=nloga all must have same bases

OpenStudy (anonymous):

log(20.25) = (log 20 + log 25) = (log (2.5.2) + log (25))= [2log(2)+ 3log(5)] 2A+ 3log(5 +1-1)=2A+3log(6/1)=2A + 3(log3 + log 2- log1)=2A+3A+3B= 5A +3B